OpenStudy (anonymous):

the measure of any angle is 12 less than twice the measure of its supplement. What is the measure of the angle?

OpenStudy (anonymous):

x = 2(180-x) - 12

OpenStudy (anonymous):

x = 114

OpenStudy (anonymous):

Sum of supplementary angles is 180 Let measure of the supplementary angle is x The given angle is 2x -12 x +2x -12 = 180 Solve for x 3x -12 = 180 3x = 192 x = 192/3 = 64

OpenStudy (anonymous):

Supplementary angle is 64, angle will be 2x -12 = 2*64 -12 = 128 -12 = 116

OpenStudy (anonymous):

Thank you. We were able to follow you on this. Thanks Rita1975 -katch